EduFarming
Your list · 0हिं

Type at least two letters. Press Esc to close.

↑↓ navigate↵ openTAB cycle filter

Conditional Probability & Bayes' Theorem · Question 88 of 125

Question 88

MediumOriginal question

P(A)=0.4P(A) = 0.4, P(B∣A)=0.7P(B \mid A) = 0.7 and P(B)=0.46P(B) = 0.46. What is P(B∣A′)P(B \mid A')?

Options for question 88

Show the answer and explanation
CorrectNot quiteAnswer B

Correct answer: B

Explanation

  1. P(B)=P(A)P(B∣A)+P(A′)P(B∣A′)P(B) = P(A)P(B \mid A) + P(A')P(B \mid A'), so 0.46=0.4×0.7+0.6x0.46 = 0.4 \times 0.7 + 0.6x.

  2. 0.6x=0.46−0.28=0.180.6x = 0.46 - 0.28 = 0.18, so x=0.3x = 0.3.

Worth noting

0.18 is P(A′∩B)P(A' \cap B); it still has to be divided by P(A′)=0.6P(A') = 0.6.

Concept
Working back through total probability
Typical time
90 s
Your time
—
The rule this testsThe theorem of total probability
Practise this chapter →

Ten questions from Conditional Probability & Bayes' Theorem, at your own pace.

Practise this chapter →

Ten questions from Conditional Probability & Bayes' Theorem, at your own pace.